Ship reports banner with links Home IODP-USIO daily/weekly reports

< Previous expedition | August >

IODP Expedition 327

Juan de Fuca Ridge-Flank Hydrogeology

Daily Science Report for 31 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: The conditioning of Hole U1362A continued until 0900 hr. The bottom-hole assembly, including a C-7 coring bit, was made up by 1230 hr. The drill string was lowered to 2650 m, the ship was maneuvered, and Hole U1362A was reentered for the seventh time at 1715 hr. The hole was again washed down to total depth and swept with mud. The rotary core barrel was deployed at 2215 hr and had cored to 3019 m (347 mbsf) by midnight.
The labs are ready to receive the first core. Outreach activities continued with an exercise on sediments and microfossils followed by making smear slides, and a talk on heat flow. Weather on site: winds 24 knots gusting to 28 knot, seas 10 feet, swells 6 ft, air temperature 13.5°C, max. heave 0.6 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 30 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Hole U1362A was reentered at 0050 hr. The top of the cement plug was tagged at 2954 m (282 m below seafloor) and the entire plug was drilled by 0730 hr (to 2981 m or 309 mbsf). The rat hole beneath the casing shoe was cleaned by 1000 hr (to 3018 m or 308 mbsf) Between 1000 hr and midnight, the hole was swept with mud, washed, and reamed to clear several spots in preparation for coring.
Core description and sampling of legacy cores continued for scientists and outreach officers on board. A meeting was held to discuss postcruise research plans with the petrology/alteration group. Weather on site: winds 24 knots gusting to 36 knots, seas 10 feet, swells 6 ft, air temperature 13.5°C, max. heave 1.4 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 29 July 2010

LOCATION: Hole U1362B (Proposed Site SR-2B; 47° 45.4997'N, 127° 45.7312'W; water depth 2672 m)
Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Hole U1362B was reentered for the fifth time at 0320 hr. Cement was tagged first at 2819 m (147 m below seafloor) and then again at 2845 m (173 mbsf), implying the cement had reached 99 m inside the casing and ~25 m up the external annulus. The cement was drilled to 2934 m (262 mbsf), or ~10 m above the 10-3/4 inch casing shoe, by 1815 hr. The drill string cleared the seafloor at 1905 hr, ending the second stage of operations at Hole U1362B. The ship was back at Hole U1362A by 2015 hr to start the third stage of operations. Rig floor maintenance was conducted until midnight.
Calibration experiments continue in the physical properties lab. Outreach activities continued with a talk on ocean crust alteration followed by core description, and the continuation of building ROVs. Weather on site: winds 25 knots gusting to 28 knots, seas 10 feet, swells 6 ft, air temperature 13.5°C, max. heave 0.8 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 28 July 2010

LOCATION: Hole U1362B (Proposed Site SR-2B; 47° 45.4997'N, 127° 45.7312'W; water depth 2672 m)

SCIENCE UPDATE: The 14-3/4 inch drill bit cleared the rig floor at 0215 hr. In preparation for running casing, 272 m of 10-3/4 inch casing was made up, followed by a swellable packer, the casing hanger, and the rest of the bottom-hole assembly (BHA). The drill string was lowered to 2684 m, the ship was maneuvered, and Hole U1362B was reentered at 1150 hr. The casing was washed to the total depth of 2944 m (272 m below seafloor) without any problems, and the casing hanger landed at 1330 hr. The hole was cemented by 1515 hr, and the drill string cleared the rig floor at 2030 hr. Because there were indications that the cement may have set prematurely too high in the casing annulus, the decision was made to verify the top of the cement plug before proceeding to Hole U1362A. The rest of the day was spent making up the 9-7/8 inch drilling BHA and running the drill string to a depth of 996 m.
A core flow and sampling meeting was held with the science party, Curator, and Staff Scientist. Calibration experiments continue in the physical properties lab. Outreach activities included a presentation on recovering CORK data with submersibles, a talk on structural geology, and the continuation of building ROVs. Weather on site: winds 20 knots gusting to 22 knots, seas 10 feet, swells 5 ft, air temperature 14°C, max. heave 0.6 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 27 July 2010

LOCATION: Hole U1362B (Proposed Site SR-2B; 47° 45.4997'N, 127° 45.7312'W; water depth 2672 m)

SCIENCE UPDATE: Drilling of the cement plug inside the 16 inch casing continued in Hole U1362B until 2916 m (244 m below seafloor). Between 0100 and 0600 hr, the hole was cleaned down to 2922 m (250 mbsf). Drilling of the 14-3/4 inch hole continued until 1645 hr to the desired depth of 2954 m (282 mbsf). The hole was conditioned for the rest of the day in preparation for running casing, and the drill string cleared the reentry cone at 2230 hr.
Outreach activities included a lecture on ocean crust petrology, a session on hard rock description, and the continuation of building ROVs. Weather on site: winds 20 knots gusting to 22 knots, seas 10 feet, swells 5 ft, air temperature 14°C, max. heave 0.6 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 26 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.6628'N, 127° 45.6720'W; water depth 2672 m)
Hole U1362B (Proposed Site SR-2B; 47° 45.4997'N, 127° 45.7312'W; water depth 2672 m)

SCIENCE UPDATE: The Hole U1362A casing was washed down to 2980.5 m (308.5 m below seafloor) and the casing hanger landed at 0530 hr. The bottom of the hole was cemented and the second stage of operations at Hole U1362A ended at 1245 hr. The second stage of operations at Hole U1362B began with the 14-3/4 inch bottom-hole assembly made up, the drill string deployed to 2655 m, and the ship maneuvered to Hole U1362B. Hole U1362B was reentered for the third time at 2030 hr, the top of cement was tagged at 2885 m (213 mbsf), and by the end of the day drilling had advanced to 2907 m (235 mbsf).
Preparation activities continue in the labs. Outreach activities included a presentation on logging tools and data, and the beginning of assembling ROVs. Weather on site: winds 22 knots gusting to 26 knots, seas 10 feet, swells 6 ft, air temperature 14°C, max. heave 1.3 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 25 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: The crew spent several hours trying to free the casing that had become stuck and by 0515 they resumed trying to wash the casing down to the desired depth of 336 m below seafloor (mbsf). Additional attempts to wash the casing past 311 mbsf failed and the decision was made at 1000 hr to shorten the casing string. After the drill string was pulled onto the rig floor, the packer was inspected at the moonpool and found ok for redeployment. The packer with short sections of casing was laid out as a unit and the next two joints of casing (27.5 m in length) were removed. The packer unit was again attached to the casing hanger and the casing was lowered to the seafloor. Hole U1362A was reentered at 2310 hr and by the end of the day the bottom of the casing had reached 2899 m (227 mbsf).
Preparation activities continue in the labs. Outreach activities included a presentation on the structure of oceanic crust. Weather on site: winds 30 knots gusting to 35 knots, seas 15 feet, swells 8 ft, air temperature 14°C, max. heave 1.3 m, max. pitch 3°, max. roll 3°.


Daily Science Report for 24 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Conditioning of Hole U1362A continued until 0200 hr, the pipe had been pulled out of the hole by 0730, and the crew prepared to run casing. 24 joints of 10 3/4 inch casing (336 m in length) were made up including a swellable packer at the top of the casing string and a casing hanger. The purpose of the packer is to isolate the borehole from the overlying ocean. Hole U1362A was reentered at 1900 hr and by the end of the day the bottom of the casing had reached 2991 m (319 m below seafloor).
Science and engineering teams continued with CORK preparations and measurements of a few Leg 168 samples using the new analytical systems. Outreach activities included a tutorial on the sample request application and a discussion of the previous day’s core observations. Weather on site: winds 30 knots gusting to 41 knots, seas 15 feet, air temperature 14.5°C, max. heave 1.5 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 23 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Drilling at Hole U1362A continued until 0230 hr to the target depth of 3018 m below rig floor (346 m below seafloor). The rest of the day was spent conditioning the hole to work through several tight spots in preparation to install the 10-3/4 inch casing.
Science and engineering teams continued with CORK preparations and measurements of a few Leg 168 samples using the new analytical systems. Outreach activities included a tutorial on the sample request application and a discussion of the previous day’s core observations. Weather on site: winds 22 knots gusting to 28 knots, seas 8 feet, swells 5 feet, air temperature 13°C, max. heave 0.8 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 22 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: The entire day was spent drilling from 2970 to 3016 m (298 to 344 m below seafloor) in basaltic oceanic crust. Mud sweeps were necessary at 2977, 2992, 3004, and 3016 m to condition the hole.
Science and engineering teams continued with CORK preparations. Outreach activities included core description and thin section sessions. Weather on site: winds 25 knots gusting to 28 knots, seas 10 feet, swells 6 feet, air temperature 13.5°C, max. heave 0.7 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 21 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Drilling the 14-3/4 inch hole for the 10 inch casing at Hole U1362A continued from 2922 to 2970 m below rig floor (250 to 298 m below seafloor). At 1545 hr, the pipe became stuck at 298 mbsf and it took an hour to work it free. The pipe was raised above the problematic interval and the hole was washed, reamed, and swept with mud over several hours to improve hole conditions. By the end of the day, the hole had been advanced another 0.5 m.
Science and engineering teams continue with CORK preparations. Outreach activities included a presentation by Geoff Wheat on building and using ROVs in schools. Weather on site: winds 22 knots gusting to 26 knots, seas 10 feet, swells 5 feet, air temperature 14°C, max. heave 1.0 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 20 July 2010

LOCATION: Hole U1362B (Proposed Site SR-2B; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)
Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: The rig floor crew continued to pull the drill string out of Hole 1362B and by 0600 hr it had cleared the rig floor. The 14-3/4 inch drilling assembly was made up and deployed through the moonpool. The vessel was offset back to Hole 1362A and the drill string reentered the hole at 1525 hr. The top of cement was encountered at 202 m below seafloor (mbsf) and drilling continued to 250 mbsf.
The physical properties team continued working with the thermal conductivity instrument and the P-wave velocity gantry. Outreach activities included a videoconference with the National Marine Educators Association (NMEA) annual conference and a discussion by Beth Orcutt on assembling colonization experiments for the microbiological studies planned for this expedition. Weather on site: winds 22 knots gusting to 26 knots, seas 10 feet, swells 5 feet, air temperature 14°C, max. heave 0.8 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 19 July 2010

LOCATION: Hole U1362B (Proposed Site SR-2B; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: The drilling assembly with the under reamer and drill bit was pulled out of the hole at 0630 hr and the rig floor crew prepared to run casing. Another sample was recovered from the drill bit and was given to the science party. The 16 inch casing was made up, the drill string was assembled, and Hole U1362B was reentered at 1500 hr. The casing was washed down to 2914 m below rig floor (242 m below seafloor) and the casing hanger was latched onto the 20 inch casing installed previously. The bottom of the hole was cemented by 2000 hr and the drill string was pulled out of the hole.
The petrologists described the ocean crust material recovered from the drill bit at Hole U1362B. The physical properties team continued working with the thermal conductivity instrument and the P-wave velocity gantry. The outreach curriculum resumed with the calibration of the tracer gas injection system. Weather on site: winds 20 knots gusting to 24 knots, seas 10 feet, swells 6 feet, air temperature 14°C, max. heave 0.6 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 18 July 2010

LOCATION: Hole U1362B (Proposed Site SR-2B; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: After verifying the depth of the reentry cone, Hole U1362B was reentered and drilling of the 21.5 inch diameter hole began. The basement contact was encountered at 242 m below seafloor (mbsf) and a smaller (18.5 inch) hole was drilled with the drill bit in advance of the under reamer to 250 mbsf. Before pulling out of the hole, mud was circulated to clean out 8 m of fill from the bottom of the hole.
Science teams handed in the first draft of their laboratory methods for review. The physical properties team practiced with the thermal conductivity instrument and the P-wave velocity gantry. Fabrication of inflatable packers for the observatories was not completed before departure from Victoria, so these were delivered to the ship by the Ocean Ranger, which arrived at 0715 hr. The transfer of cargo and three Transocean staff was completed by 0920 hr. Weather on site: winds 25 knots gusting to 26 knots, seas 6 feet, swells 8 feet, air temperature 14°C, max. heave 0.5 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 17 July 2010

LOCATION: Hole U1362B (Proposed Site SR-2B; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: Installation of the Hole U1362B reentry cone and casing was completed with the casing jetted in to a total depth of 53 meters below seafloor. After the drill string was pulled out of the hole, the bottom-hole assembly with a drill bit and under reamer was assembled, and the drill string was lowered to the seafloor. The under reamer is used to create a 21 inch hole for the 16 inch casing.
The petrologists continued working on the ocean crust material recovered from the drill bit at Hole U1362B. The outreach curriculum resumed with scientific illustration in the afternoon. Marine lab specialists have replaced the carpet with wood flooring in the internet cafe and are working on a variety of projects in the labs including refurbishing the Carver presses in the Chemistry lab and the Section Half Multisensor Logger in the Core lab. Weather on site: winds 20 knots gusting to 22 knots, seas 10-12 feet, swells 6 feet, air temperature 14°C, max. heave 0.6 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 16 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2A; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)
Hole U1362B (Proposed Site SR-2B; 47° 45.499'N, 127° 45.733'W; water depth 2672 m)

SCIENCE UPDATE: The casing at the bottom of Hole U1362A was cemented and the vessel was offset to the coordinates of Hole U1362B. The reentry cone, running tool, and 20 inch casing were assembled and deployed through the moonpool. Hole U1362B was spudded at 2135 hr and the reentry cone and casing were jetted into 21 meters below seafloor.
CORK specialists continued with their preparations. The petrologists worked on the core description template and started describing ocean crust material recovered from the drill bit. The outreach curriculum resumed in the morning with a discussion on the Darcy lab and a reading assignment for the Journal Club on How People Learn. Weather on site: winds 23 knots gusting to 25 knots, seas 12 feet, swells 10 feet, air temperature 13.5°C, max. heave 0.6 m, max. pitch 2°, max. roll 2°.


Daily Science Report for 15 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: The debris at the bottom of Hole U1362A was cleaned by circulating high-viscosity mud, the under reamer and bit were brought up to the rig floor at 0920 hr, 230 meters of 16 inch casing was assembled, the hole was reentered at 2015 hr, and the casing was ran into the hole and had reached the target of 230 mbsf by 2315 hr. An operations meeting was held to discuss cementing.
A science meeting was held on the description and sampling of upper oceanic crust material recovered from the drill bit. CORK specialists continued with their preparations. The outreach curriculum resumed in the morning with part 2 of the Darcy lab, a presentation on IODP databases, and training on recording audio podcasts. Weather on site: winds 22 knots gusting to 26 knots, seas 12 feet, swells 9 feet, air temperature 14°C, max. heave 1.0 m, max. pitch 2°, max. roll 3°. Correction to yesterday's report: basement contact placed at 236 mbsf.


Daily Science Report for 14 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Hole U1362A was reentered at 0245 hr with the drilling assembly and drilling continued from 53 to 241 meters below seafloor (mbsf). The sediment/basement contact has been placed at 232 mbsf. Operations, engineering, and scientific personnel met to discuss the CORK and instrument string deployment plans for Sites U1362, U1301, and 1027.
The petrologists continued describing legacy core from Leg 168, Site 1027. CORK specialists continued with their preparations. The outreach curriculum resumed in the morning with an introduction to Osmosamplers and a series of experiments that will take place over the next several days. Dinah Bowman presented a class on pencil drawing in the afternoon. Weather on site: winds 18 knots gusting to 24 knots, seas 4 feet, swells 6 feet, air temperature 14°C, max. heave 0.6 m, max. pitch 2°, max. roll 3°.


Daily Science Report for 13 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: The Hole U1362A reentry cone and 20 inch casing were jetted in to 50.5 meters below seafloor (52.5 mbsf when corrected for the daily tide). The cone was placed perfectly with its base plate on the seafloor. The cone was released from the running tool, the subsea TV was recovered, and the drill string was pulled out of the hole by 1800. A drilling assembly was made up and had reached 1863 meters below rig floor by the end of the day.
The petrologists have started describing legacy core from Leg 168, Site 1027. The outreach officers were trained on the digital imaging track and commenced imaging Leg 168 cores. CORK specialists continued with their preparations. The outreach curriculum continued with a multiday Darcy's Law experiment led by Andy Fisher and Dustin Winslow. Weather on site: winds 18 knots gusting to 22 knots, seas 10 feet, swells 8 feet, air temperature 14°C, max. heave 1.5 m, max. pitch 2°, max. roll 3°.


Daily Science Report for 12 July 2010

LOCATION: Hole U1362A (Proposed Site SR-2; 47° 45.662'N, 127° 45.674'W; water depth 2672 m)

SCIENCE UPDATE: Completed jet-in test at Hole U1362A, made up reentry cone/casing running tool, assembled 20 inch casing, deployed reentry cone/casing through moonpool, spudded Hole U1362A at 2345 hr, and proceeded to jet in reentry cone with 53 meters of casing from the seafloor to 12 meters below seafloor.
Training continued in the core description and physical properties labs. CORK specialists continued with their preparations. All scientists and outreach officers participated in a sampling policy, SampleMaster, and core flow orientation. Stefan Mrozewski resumed the outreach curriculum with "Drilling, Coring, Logging, and CORKing." Weather on site: winds 25 knots gusting to 30 knots, seas 9 feet, swells 14 feet, air temperature 14°C, max. heave 5 feet, max. pitch 2°, max. roll 2°.


Daily Science Report for 11 July 2010

LOCATION: Site U1362 (Proposed Site SR-2; 47° 45.5'N, 127° 45.7'W; water depth 2672 m)

SCIENCE UPDATE: Completed jet-in test at Hole U1362B, offset ship ~300 m in DP mode, and initiated jet-in test at Hole U1362A.
Training continued in the core description and physical properties labs. CORK specialists continued with their preparations. Following the World Cup final, Leslie Peart kicked off the outreach curriculum with "Lecture as a Last Resort." Weather on site: winds 20 knots gusting to 30 knots, seas 5 feet, swells 9 feet, air temperature 14°C, heave 1 foot, pitch 0.5°, roll 1°.


Daily Science Report for 10 July 2010

LOCATION: Arrived at Site U1362 (Proposed Site SR-2; 47° 45.5'N, 127° 45.7'W)

SCIENCE UPDATE: Arrived on site at 0815 hr and deployed positioning beacon at 0922 hr. Began assembling bottom hole assembly for jet-in test at Hole U1362B. Seafloor was tagged at 2672 meters below rig floor (mbrf) at 2200 hr.
Training continued in the core description and physical properties labs. Engineers finished plumbing the CORK instrument bays and the CORK specialists continued with their preparations. Weather on site: winds 20 knots, 6 foot swells, air temperature 14°C, heave 0.7 feet, 1° pitch, 1° roll.


Daily Science Report for 9 July 2010

LOCATION: En route to Prospectus Site SR-2 (Site U1362)

SCIENCE UPDATE: Completed port call loading activities and secured cargo for transit. Last line away at 12:12 pm. Scientists continue to organize their lab areas. Petrologists started training on core description software. Engineers continued plumbing the instrument bays. Weather en route: 28 knot winds gusting to 32 knots, 5 foot seas/swells, 1° pitch, 2° roll. Anticipated arrival on site at 8 am tomorrow.


Daily Science Report for 8 July 2010

LOCATION: Esquimalt Graving Dock, Victoria, British Columbia, Canada

SCIENCE UPDATE: Meetings continued with the Captain's safety orientation, engineering update, lab space allocation, and outreach goals and activities. The outreach officers received training on the joidesresolution.org web site. Successfully completed PHC high pressure testing. Technical staff continued to prepare the labs. The frozen food shipment was loaded onboard. Loading of fuel was completed in the afternoon. Departure is scheduled for noon on July 9.


Daily Science Report for 7 July 2010

LOCATION: Esquimalt Graving Dock, Victoria, British Columbia, Canada

SCIENCE UPDATE: Orientations started with scientist, outreach, and technical staff introductions. Presentations included expedition science objectives by Co-Chief Scientist Andy Fisher, life on board, what we are going to accomplish during the expedition and how we will do it, and IT resources. Scientists and outreach officers participated in lab safety tours. Visiting scientists from the Pacific Geoscience Centre continued to work on the CORK pressure instrument bays. The reentry cone for Site SR-2A was assembled. Technical staff continued to load supplies and prepare the labs. Perforated drill collars and CORK running tools arrived and were loaded onboard, as well as dry goods and refrigerated food. Loading of fuel started in the afternoon. Last two scientists moved onboard. SCIMPI personnel (Moran, Farrington, Newman, and Kulin) toured the JR and met with USIO and ODL management and engineering personnel. Departure is scheduled for the morning of July 9.


Daily Science Report for 6 July 2010

LOCATION: Esquimalt Graving Dock, Victoria, British Columbia, Canada

SCIENCE UPDATE: Science and outreach participants arrived. Scientists started unpacking their scientific equipment. Visiting scientists from the Pacific Geoscience Centre worked on the CORK pressure instrument bays. The education officers participated in Deep Earth Academy orientations. The group included the education officers for the South Pacific Gyre and Mid-Atlantic Microbiology expeditions. Heavy cargo was secured on deck. Technical staff continued to load supplies and prepare the labs.


Daily Science Report for 5 July 2010

LOCATION: Esquimalt Graving Dock, Victoria, British Columbia, Canada

SCIENCE UPDATE: Completed USIO crew change and crossover. TAMU surface freight offloaded. CORKs arrived and were loaded onboard.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Tuesday, 08-Nov-2011 08:50:38 CST.